### v3

@@ Abstract
Railway networks have become increasingly important in recent times, especially to move freight and public transportation from road traffic and planes to more environmentally friendly trains. Since expanding the global railway network is time and resource-consuming, maximizing the rail capacity on the existing infrastructure is desirable. However, simply running more trains is infeasible as certain constraints enforced by the train control system must be satisfied. The capacity of a network depends (amongst others) on the distance between trains allowed by this safety system. While most signaling systems rely on fixed blocks defined by costly hardware, new specifications provided by Level 2 with Hybrid Train Detection of the European Train Control System (ETCS L2 HTD), formerly known as ETCS Hybrid Level 3, allow the usage of virtual subsections. This additional degree of freedom allows for shorter train following times and, thus, more trains on existing railway tracks. On the other hand, new design tasks arise on which automated methods might be helpful for designers of modern railway networks. However, although first approaches exist that solve design problems arising within ETCS L2 HTD, neither formal descriptions nor results on the computational complexity of the corresponding design tasks exist. In this paper, we fill this gap by providing a formal description of design tasks for ETCS L2 HTD and proof that these tasks are NP-complete or NP-hard, respectively. By that, we are providing a solid basis for the future development of methods to solve those tasks, which will be integrated into the Munich Train Control Toolkit available open-source on GitHub at https://github.com/cda-tum/mtct.
@@ I-B Related Work
This additional degree of freedom allows for shorter train following times and, thus, more trains on existing railway tracks. At the same time, new design tasks arise that require automated methods to be helpful for designers of modern railway networks. However, thus far, design automation methods for ETCS, on the one hand, mainly focus on blocks characterized by hardware (e.g., axle counters). Because of this, they use general measures of capacity, energy efficiency, or economic efficiency independently from a specific schedule to propose a signaling layout [14, 15, 16, 17, 18, 19, 20, 21, 22]. Thus, the respective solution is expected to work well on various schedules. On the other hand, train routing was considered separately, assuming the block layout had already been decided upon. Approaches are mainly separated in macroscopic routing [23] and track allocation on a microscopic level, i.e., considering the constraints imposed by the control system and sections. Track allocation was considered in general [24, 25, 26] and specifically focusing on railway stations [27, 28, 29]. Besides that, recent research by Schlechte et al. optimizes time tables on dedicated moving block systems [30], which was later extended to use lazy constraints to speed up computation [31, 32]. However, implementing a moving block system in practice is significantly harder than using virtual subsections, in which case only the design but not the control logic changes [12].
@@
While (almost) all of these methods can also be applied for ETCS L2 HTD, because virtual subsections and TTD sections have essentially the same consequences in final operation (assuming no system fails), they are not utilizing the additional degree of freedom introduced by virtual subsections. Using ETCS L2 HTD, changing a previous block layout defined by TTD sections even after building the network is now possible. In particular, it might be reasonable to consider a desired schedule directly because changing a block layout with a new schedule is more manageable than before. However, corresponding considerations are just at the beginning.
@@
For ETCS L2 HTD, specifically, the contributions to the “Hybrid ERTMS/ETCS Level 3 Case Study” provided an step in this direction [33]. Here, the main goal was to detect potential ambiguities in the specification to fix them in later versions. Submitted papers do this using iUML-B, Event-B, Formal B, Electrum, SysML/KAOS, and SPIN [33]. However, while these formalizations provided initial (non-ambiguous) interpretations of the ETCS L2 HTD, no precise design tasks, e.g., how to generate new layouts, can be derived from that. To the best of our knowledge, only research proposed in [34, 35, 10, 36] explicitly considered the automatic design of hybrid train detection systems. However, even those papers specified the design tasks at a relatively high level, and neither formal descriptions nor results on the computational complexity of those tasks have been provided.
@@
Overall, all related work on developing design methods for ETCS L2 HTD has been conducted without a clear definition nor a formal understanding of the complexity of the underlying task. As a consequence, we currently have heuristics that usually generate non-optimal results, on the one hand, [35] and exact solutions which significantly suffer from huge run-times and severely limited scalability on the other [34, 10, 36], but do not know whether this is caused by insufficient methodology or complexity reasons. Similar scheduling problems (e.g., [24]) have been proven to be NP-hard by using the alternative graph introduced in [37]. At the same time, we are not aware of formal complexity results on the design tasks considered in [34, 35, 10, 36].
@@ II-A Basic Definition of Railway Networks
Figure 3: Siding as a graph
Figure 4: Simple turnout
@@
The set of border vertices is denoted by $\mathcal{B}\subseteq V$ and all $v\in\mathcal{B}$ fulfill $|\Gamma(v)|<=1$, where $\Gamma(\cdot)$ denotes the neighboring vertices, i.e., they cannot be in the interior of the network. It is assumed that trains can only enter and exit the network with predefined headway times $h\colon\mathcal{B}\to\operatorname{\mathbb{R}}_{\geq 0}$.
@@ II-B Trains and Their Schedules
A train $\operatorname{tr}:=\left(l^{(\operatorname{tr})},v_{max}^{(\operatorname{tr})},a^{(\operatorname{tr})},d^{(\operatorname{tr})}\right)$ is a tuple, where
@@
  • $v_{max}^{(\operatorname{tr})}$ defines its maximal possible speed, and
  • $a^{(\operatorname{tr})}$ and $d^{(\operatorname{tr})}$ denote the maximal acceleration and deceleration respectively.
@@
Definition II.11 (Train Route).
For a given train $\operatorname{tr}$, a train route $R^{(\operatorname{tr})}_{[a,b]}:=\left(\operatorname{rg}_{t}^{(\operatorname{tr})}\right)_{t\in[a,b]}$ is a series of track ranges for every time $t\in[a,b]$, such that $\operatorname{len}\left(\operatorname{rg}_{t}^{(\operatorname{tr})}\right)=l^{(\operatorname{tr})}$.
@@
to be the occupied track segments. Naturally, $\operatorname{occ}_{t}^{(\operatorname{tr})}:=\operatorname{occ}\left(\operatorname{rg}_{t}^{(\operatorname{tr})}\right)$ corresponds to the segments occupied by train $tr$ at time $t\in[a,b]$.
@@
to be their distance. Using this, we can characterize the speed of a train by the following definition.
@@
Of course, further details of train movements can be modeled within this framework as additional constraints. For example, the train’s acceleration and deceleration properties might depend on its current speed and/or the underlying track’s gradient. Hence, braking curves defined in the official specifications can also be considered [41].
@@
Consider Figure 10 with an 360m-long train positioned at $(80,(e_{1},0,1),(e_{2},0,0.6),0)$, which moved to $(0,(e_{2},0.2,1),(e_{3},0,0.4))$ within $5$ seconds. Then, the distance $\operatorname{dist}\left(\operatorname{rg}_{0s}^{(tr)},\operatorname{rg}_{5s}^{(tr)}\right)$ is given by $\underbrace{(80m-0m)}_{s_{in}^{(0s)}-s_{in}^{(5s)}}+\underbrace{(1-0)}_{\mu_{1}^{(0s)}-\lambda_{1}^{(0s)}}\cdot\underbrace{100m}_{\operatorname{len}(e_{1})}+\underbrace{(0.2-0)}_{\lambda_{1}^{(5s)}-\lambda_{2}^{(0s)}}\cdot\underbrace{300m}_{\operatorname{len}(e_{2})}=240m$. Hence, we can approximate its speed by $\frac{240m}{5s}=48\frac{m}{s}$.
@@
Definition II.15 (Train Schedule).
Let $tr_{1},\ldots,tr_{m}$ be $m$ trains on a railway network $N$. For every train $tr_{i}$, a train schedule consists of
@@
A train $tr_{i}$ respects its schedule if,
@@ II-C Sections and Virtual Subsections in Railway Networks
A key concept in train control is the division of the tracks into multiple blocks. A block is only allowed to be occupied by at most one train. In the context of ETCS L2 HTD, these blocks are given by the VSS sections that function essentially analog to their TTD counterparts. Hence, the respective control condition can be formalized as follows and depicted in Figure 12.
@@
Overall, given a railway network $N$, a certain number of trains, and a train schedule, the following properties are desired whenever considering a design task:
  • Every train respects its maximal speed (Definition II.12) and schedule (Definition II.15).
@@
Figure 12: VSS condition
Again, note that one can, in theory, model the behavior of the control system in as much detail as desired. For example, one should consider that the entire braking distance in front of a train should not be occupied in the VSS condition (Definition II.17). This could, e.g., be achieved by including this in the train’s occupation as in [10]. Moreover, interlockings could be considered using additional constraints. Nevertheless, we believe the level of detail provided in this paper suffices for defining relevant design tasks.
@@ III Design Tasks and their Complexity
In this section, we define different design tasks that might occur in ETCS L2 HTD systems and could be valuable to automate. They are formulated independently from whichever solution method might be chosen to solve them.
III-A Verification of Train Schedules on ETCS L2 HTD Layouts
Note that a desired train schedule might be impossible on a particular railway network since only one train can be located on each TTD/VSS section, a restricted track speed, or a limited train speed.
@@ III-A Verification of Train Schedules on ETCS L2 HTD Layouts
Consider the train network in Figure 13 and three trains $\operatorname{tr}_{1}$ to $\operatorname{tr}_{3}$. As before, corresponds to a TTD border and to no border. Also, we are given one real-world train station $S:=\{\tilde{e_{5}},\tilde{e_{10}}\}$, which is marked in blue. The schedules are specified as follows:
@@
where time is measured in seconds. Observe that the station $S$ only consists of two blocks; hence, a maximum of two trains can stop at a time. However, the schedule requires all three trains to be stopped at the station at time $t=270$, for example. Hence, there are no train routes such that the VSS condition is fulfilled, the trains respect their schedule, and the verification task returns infeasibility.
@@
Observe that the routing on a fixed VSS layout might differ slightly from systems using trackside-oriented train detection. This is because trains equipped with train integrity monitoring (TIM) can report VSS sections free independently of their corresponding TTD sections. Any train can follow closely. At the same time, trains without TIM need TTD hardware (e.g., axle counters) to report a section as free. Hence, a train can only follow if a non-TIM train has cleared an entire TTD section.
If all trains are equipped with TIM, the situation simplifies. In that case, there is no logical difference between TTD and VSS sections, and the block layout is fixed for the verification tasks defined above. Hence, TS-VER is essentially equivalent to the task on classical layouts in this (simplified) case. Scheduling under such classical signaling systems is known to be NP-hard in general [37, 24]. In particular, it is reasonable to assume that TS-VER has similar complexity. At the same time, there are minor differences to the scheduling problems considered so that the proof does not carry over formally. While the following complexity result is not unexpected, we formally include them for reasons of completeness. Moreover, it serves as a basis for the following design tasks in Sections III-B and III-C, which are tailored to ETCS L2 HTD.
@@
TS-VER is NP-complete even if all routes and exact stopping times in train stations are fixed and the railway network is planar.
@@
Given a certificate, we can check feasibility using the definitions in polynomial time; hence, TS-VER is in NP. To show NP-completeness, we reduce Monotone 3-SAT-(2,2$\leq$ 2$\leq$ 2≤ 2 , ≤ 2) to TS-VER. Monotone 3-SAT-(2,2$\leq$ 2$\leq$ 2≤ 2 , ≤ 2) are all SAT instances in which every clause contains three literals (all of which contain only unnegated or only negated variables), and every variable appears in clauses at most twice unnegated and twice negated. Monotone 3-SAT-(2,2$\leq$ 2$\leq$ 2≤ 2 , ≤ 2) is known to be NP-complete [42].
@@
Figure 14: Reduction from Monotone 3-SAT-(2,2$\leq$ 2$\leq$ 2≤ 2 , ≤ 2) to TS-VER

### v5

@@ Abstract
Railway networks have become increasingly important in recent times, especially in moving freight and public transportation from road traffic and planes to more environmentally friendly trains. Since expanding the global railway network is time- and resource-consuming, maximizing the rail capacity of the existing infrastructure is desirable. However, simply running more trains is infeasible as certain constraints enforced by the train control system must be satisfied. The capacity of a network depends (amongst others) on the distance between trains allowed by this safety system. While most signaling systems rely on fixed blocks defined by costly hardware, new specifications provided by Level 2 with Hybrid Train Detection of the European Train Control System (ETCS L2 HTD), formerly known as ETCS Hybrid Level 3, allow the usage of virtual subsections. This additional degree of freedom allows for shorter train following times and, thus, more trains on existing railway tracks.
On the other hand, new design tasks arise on which automated methods might be helpful for designers of modern railway networks. However, although first approaches exist that solve design problems arising within ETCS L2 HTD, neither formal descriptions nor results on the computational complexity of the corresponding design tasks exist. In this paper, we fill this gap by providing a formal description of design tasks for ETCS L2 HTD and proof that these tasks are NP-complete or NP-hard, respectively. By that, we are providing a solid basis for the future development of methods to solve those tasks, which will be integrated into the Munich Train Control Toolkit available open-source on GitHub at https://github.com/cda-tum/mtct.
@@ I-B Related Work
This additional degree of freedom allows for shorter train following times and, thus, more trains on existing railway tracks. At the same time, new design tasks arise that require automated methods to be helpful for designers of modern railway networks. However, thus far, design automation methods for ETCS, on the one hand, mainly focus on blocks characterized by hardware (e.g., axle counters). Because of this, they use general measures of capacity, energy efficiency, or economic efficiency independently from a specific schedule to propose a signaling layout [14, 15, 16, 17, 18, 19, 20, 21, 22]. Thus, the respective solution is expected to work well on various schedules. On the other hand, train routing was considered separately, assuming the block layout had already been decided upon. Approaches are mainly separated in macroscopic routing [23] and track allocation on a microscopic level, i.e., considering the constraints imposed by the control system and sections. Track allocation was considered in general [24, 25, 26] and specifically focusing on railway stations [27, 28, 29].
This research has been extended to future signaling systems, such as ETCS L2 HTD and Moving Block control systems. Headways on such railway lines are different to former theory leading to adaptions in the blocking-time theory [30]. Besides that, recent research by Schlechte et al. optimizes time tables on dedicated moving block systems [31], which was later extended to use lazy constraints to speed up computation [32, 33]. However, implementing a moving block system in practice is significantly harder than using virtual subsections, in which case only the design but not the control logic changes [12].
@@
While (almost) all of these methods can also be applied for ETCS L2 HTD, because virtual subsections and TTD sections have essentially the same consequences in final operation (assuming no system fails), they are not utilizing the additional degree of freedom introduced by virtual subsections. Using ETCS L2 HTD, changing a previous block layout defined by TTD sections even after building the network is now possible. In particular, it might be reasonable to consider a desired schedule directly because changing a block layout with a new schedule is more manageable than before. However, corresponding joint considerations of infrastructure planning and scheduling are just at the beginning.
@@
For ETCS L2 HTD, specifically, the contributions to the “Hybrid ERTMS/ETCS Level 3 Case Study” provided an step in this direction [34]. Here, the main goal was to detect potential ambiguities in the specification to fix them in later versions. Submitted papers do this using iUML-B, Event-B, Formal B, Electrum, SysML/KAOS, and SPIN [34]. However, while these formalizations provided initial (non-ambiguous) interpretations of the ETCS L2 HTD, no precise design tasks, e.g., how to generate new layouts, can be derived from that. To the best of our knowledge, only research proposed in [35, 36, 10, 37] explicitly considered the automatic design of hybrid train detection systems based on specific timetable requests. However, even those papers specified the design tasks at a relatively high level, and neither formal descriptions nor results on the computational complexity of those tasks have been provided.
@@
Overall, all related work on developing design methods for ETCS L2 HTD has been conducted without a clear definition nor a formal understanding of the complexity of the underlying task. As a consequence, we currently have heuristics that usually generate non-optimal111with respect to some objective describing the operational outcome, see also Section III results, on the one hand, [36] and exact solutions which significantly suffer from huge run-times and severely limited scalability on the other [35, 10, 37], but do not know whether this is caused by insufficient methodology or complexity reasons. Similar scheduling problems (e.g., [24]) have been proven to be NP-hard by using the alternative graph introduced in [38]. At the same time, we are not aware of formal complexity results on the design tasks considered in [35, 36, 10, 37].
@@ II-A Basic Definition of Railway Networks
Figure 3: Siding as a graph
Figure 4: Simple turnout
@@
The set of border vertices is denoted by $\mathcal{B}\subseteq V$ and all $v\in\mathcal{B}$ fulfill $|\Gamma(v)|=1$, where $\Gamma(\cdot)$ denotes the neighboring vertices, i.e., they cannot be in the interior of the network. It is assumed that trains can only enter and exit the network with predefined headway times $h\colon\mathcal{B}\to\operatorname{\mathbb{R}}_{\geq 0}$.
@@ II-B Trains and Their Schedules
A train $\operatorname{tr}:=\left(l^{(\operatorname{tr})},v_{max}^{(\operatorname{tr})},a^{(\operatorname{tr})},d^{(\operatorname{tr})},\operatorname{tim}^{(\operatorname{tr})}\right)$ is a tuple, where
@@
  • $v_{max}^{(\operatorname{tr})}$ defines its maximal possible speed,
  • $a^{(\operatorname{tr})}$ and $d^{(\operatorname{tr})}$ denote the maximal acceleration and deceleration respectively, and
  • $\operatorname{tim}^{(\operatorname{tr})}\in\{true,false\}$ denotes if $\operatorname{tr}$ is equipped with train integrity monitoring (TIM).
@@
At this point, it would be intuitive to define the position of a train at a given time as a track range $\operatorname{rg}$ with $\operatorname{len}(\operatorname{rg})=l^{(\operatorname{tr})}$. However, it is convenient to use a slightly different definition, which later makes it easier to properly define separation trains (or headways) of different trains with respect to each other. Note that the entire distance needed to brake has to be already cleared from other trains. Hence, one could say that a train does not only occupy the track parts on which it is physically present, but also all of its braking distance ahead.
Definition II.11 (Train Route and Occupation).
For a given train $\operatorname{tr}$, a train route $R^{(\operatorname{tr})}_{[a,b]}:=\left(\operatorname{rg}_{t}^{(\operatorname{tr})}\right)_{t\in[a,b]}$ is a series of track ranges for every time $t\in[a,b]$, such that $\operatorname{len}\left(\operatorname{rg}_{t}^{(\operatorname{tr})}\right)=l^{(\operatorname{tr})}+\operatorname{bd}_{t}^{(\operatorname{tr})}$, where $\operatorname{bd}_{t}^{(\operatorname{tr})}$ denotes the distance $\operatorname{tr}$ requires to come to a full stop if it were to start decelerating at time $t$. In this context $\operatorname{rg}_{t}^{(\operatorname{tr})}$ is bound by the trains rear end on the one and its moving authority on the other side.
@@
to be the occupied track segments. Naturally
$\operatorname{occ}_{t}^{(\operatorname{tr})}:=\operatorname{occ}\left(\operatorname{rg}_{t}^{(\operatorname{tr})}\right)$ (12)
corresponds to the segments occupied by train $tr$ at time $t\in[a,b]$.
@@
to be their distance. Using this, we can characterize the varying speed of a train at time $t$ by the following definition.
@@
A train respects its braking distance if for every $t\geq 0$
$\operatorname{bd}_{t}^{(\operatorname{tr})}\geq\frac{v_{t}^{(\operatorname{tr})}}{2\cdot d^{(\operatorname{tr})}}.$ (17)
By definition, $t\in\mathbb{R}$, so that train velocity and occupation are defined continuously. When solving design tasks proposed in Section III, time might or might not have to be discretized depending on the applied algorithm. However, this does not affect the problem description itself, which assumes continuous time and is independent of proposed solving strategies.
Of course, further details of train movements can be modeled within this framework as additional constraints. For example, the train’s acceleration and deceleration properties might depend on its current speed and/or the underlying track’s gradient. Hence, braking curves defined in the official specifications can also be considered [42] by altering Equations (16) and (17) respectively. Usually, trains should not face red signals when planning their movements. Such a requirement could, e.g., be included as further constraints on the speed to disallow stopping on an open line.
@@
Consider Figure 10 with an 210m-long train occupying $(80,(e_{1},0,1),(e_{2},0,0.6),0)$, which moved to $(0,(e_{2},0.2,1),(e_{3},0,0.4))$ within $10$ seconds (we have $\operatorname{bd}_{0s}^{(\operatorname{tr})}=\operatorname{bd}_{10s}^{(\operatorname{tr})}=150m$). The distance $\operatorname{dist}\left(\operatorname{rg}_{0s}^{(\operatorname{tr})},\operatorname{rg}_{10s}^{(\operatorname{tr})}\right)$ is given by $\underbrace{(80m-0m)}_{s_{in}^{(0s)}-s_{in}^{(10s)}}+\underbrace{(1-0)}_{\mu_{1}^{(0s)}-\lambda_{1}^{(0s)}}\cdot\underbrace{100m}_{\operatorname{len}(e_{1})}+\underbrace{(0.2-0)}_{\lambda_{1}^{(10s)}-\lambda_{2}^{(0s)}}\cdot\underbrace{300m}_{\operatorname{len}(e_{2})}=240m$. Hence, we can approximate its speed by $\frac{240m}{10s}=24\frac{m}{s}$.
@@
Definition II.15 (Timetable Requests).
Let $tr_{1},\ldots,tr_{m}$ be $m$ trains on a railway network $N$. For every train $tr_{i}$, a timetable request consists of
@@
A train $tr_{i}$ respects its timetable request if,
@@ II-C Sections and Virtual Subsections in Railway Networks
A key concept in train control is the division of the tracks into multiple blocks. A block is only allowed to be occupied by at most one train. In the context of ETCS L2 HTD, these blocks are usually given by the VSS sections that function essentially analog to their TTD counterparts.
The concrete control logic depends on if a train is equipped with train integrity monitoring (TIM) or not. A train with TIM can safely reports its rear end, because it is determined with sufficient certainty that the train is still complete and has not lost any wagons. Because of this a VSS section can be safely reported as free without the need of any TTD hardware. Hence, any other train is permitted to enter this VSS that has just been cleared. On the other hand, a train without TIM can only safely report its front position. Thus, a VSS remains occupied until the corresponding TTD section is cleared by the trackside hardware. These cases are depicted in Figure 12, in which only 4 out of 6 trains are equipped with TIM.
Figure 12: VSS condition
Hence, the respective control condition can be formalized as follows:
@@
Furthermore, let $T_{1},\ldots,T_{l}$ be the TTD-sections as defined in Equation (19). For any $h\in\{1,\ldots,l\}$, $j\in\{1,\ldots,m\}$, and $t\geq 0$ with $T_{h}\cap\operatorname{occ}_{t}^{(\operatorname{tr}_{j})}\neq\emptyset$ let
$\hat{t}^{(\operatorname{tr}_{j})}_{h}(t):=\max\left\{0,\sup\left\{\tau\colon 0\leq\tau\leq t,T_{h}\cap\operatorname{occ}_{\tau}^{(\operatorname{tr}_{j})}=\emptyset\right\}\right\}.$ (22)
Using the above definition, for every $j\in\{1,\ldots,m\}$ with $\operatorname{tim}^{(\operatorname{tr}_{j})}=false$ as well as $k\in\{1,\ldots,m\}-\{j\}$, $h\in\{1,\ldots,l\}$, $i\in\{1,\ldots,n\}$, $t\geq 0$ and $\tau\in\left[\hat{t}^{(\operatorname{tr}_{j})}_{h}(t),t\right]$ it has to additionally hold that
$T_{h}\cap V_{i}\cap\operatorname{occ}_{\tau}^{(\operatorname{tr}_{j})}\neq\emptyset\Rightarrow V_{i}\cap\operatorname{occ}_{t}^{(\operatorname{tr}_{k})}=\emptyset.$ (23)
Intuitively, in Definition II.17, Equation (21) ensures that only one VSS is occupied at any given time. $\hat{t}^{(\operatorname{tr}_{j})}_{h}(t)$ is the time at which $\operatorname{tr}_{j}$ entered (began to occupy) $T_{h}$. Equation (23) simply states that if a train without TIM occupies a TTD section $T_{h}$ any VSS within $T_{h}$ previously occupied by that train is not yet cleared.
Because the braking distance was already included in each trains occupation (see Definition II.11), the conditions in Definition II.17 ensure the safe distance required between following trains.
Remark II.18.
If all trains are equipped with TIM, only Equation (21) remains in Definition II.17. Using an ETCS L2 control with fixed TTD sections and no VSS, Definition II.17, again, simplifies to Equation (21) with $T_{1},\ldots,T_{h}$ (TTD-sections) instead of $V_{1},\ldots,V_{n}$ (VSS-sections), i.e., $T_{i}\cap\operatorname{occ}_{t}^{\left(tr_{j}\right)}\neq\emptyset\Rightarrow T_{i}\cap\operatorname{occ}_{t}^{\left(tr_{k}\right)}=\emptyset$.
Overall, given a railway network $N$, a certain number of trains, and a list of timetable requests, the following properties are desired whenever considering a design task:
  • Every train respects its maximal speed and braking distance (Definition II.12)
  • Every train respects its timetable request (Definition II.15).
@@
Again, note that one can, in theory, model the behavior of the control system in as much detail as desired. For example, interlockings or time needed for turnouts to change their position could be considered using additional constraints. Nevertheless, we believe the level of detail provided in this paper suffices for defining relevant design tasks.
@@ III Design Tasks and their Complexity
In this section, we define different design tasks that might occur in ETCS L2 HTD systems and could be valuable to automate. They are formulated independently from whichever solution method might be chosen to solve them. In Section III-A, we formulate a respective verification tasks. While the verification is arguably more of a theoretical nature, it provides the basis for the design tasks proposed in Sections III-B and III-C. Especially the optimization tasks in Section III-C, which use the added flexibility by HTD control systems to use specific timetable requests instead of general schedule-independent KPIs, are both, of practical interest and not yet well studied.
The proposed design tasks in the following subsections are formulated to help infrastructure planners to find optimal signaling layouts in terms of enabling the best operational outcome based on specific timetable requests. They do not yet consider robustness with respect to delays. For classical signaling systems capacity was already considered under disturbed conditions leading to real-time rescheduling problems [43]. Recently such rescheduling problems have also been considered for modern (moving block) control systems [44]. At the same time, previous work focuses on minimizing the effect of delays on a fixed signaling layout. An interesting extension to the presented design tasks would be to also consider stochastic distributions on expected delays. In that case one could aim to optimize the VSS layout to minimize the negative effects of expected rescheduling during operation. At the time of writing, the proposed design tasks are not well studied even in the deterministic case. Hence, for now, we focus on providing efficient solutions for the proposed design tasks. Once reasonable methods exist, adding expected stochastic delays would be a natural extension.
III-A Verification of Timetable Requests on ETCS L2 HTD Layouts
Note that a desired timetable request might be impossible on a particular railway network since only one train can be located on each TTD/VSS section, a restricted track speed, or a limited train speed.
@@ III-A Verification of Timetable Requests on ETCS L2 HTD Layouts
Consider the train network in Figure 13 and three trains $\operatorname{tr}_{1}$ to $\operatorname{tr}_{3}$. As before, corresponds to a TTD border and to no border. Also, we are given one real-world train station $S:=\{\tilde{e_{5}},\tilde{e_{10}}\}$, which is marked in blue. The timetable requests are specified as follows:
@@
where time is measured in seconds. Observe that the station $S$ only consists of two blocks; hence, a maximum of two trains can stop at a time. However, the timetable requests require all three trains to be stopped at the station at time $t=270$, for example. Hence, there are no train routes such that the VSS condition is fulfilled, the trains respect their timetable requests, and the verification task returns infeasibility.
@@
Observe that the routing on a fixed VSS layout might differ slightly from systems using trackside-oriented train detection, see also discussion around Definition II.17. This is because trains equipped with train integrity monitoring (TIM) can report VSS sections free independently of their corresponding TTD sections. Any train can follow closely. At the same time, trains without TIM need TTD hardware (e.g., axle counters) to report a section as free. Hence, a train can only follow if a non-TIM train has cleared an entire TTD section.
If all trains are equipped with TIM, the situation simplifies. In that case, there is no logical difference between TTD and VSS sections, and the block layout is fixed for the verification tasks defined above. Hence, TR-VER is essentially equivalent to the task on classical layouts in this (simplified) case. Scheduling under such classical signaling systems is known to be NP-hard in general [38, 24]. In particular, it is reasonable to assume that TR-VER has similar complexity. At the same time, there are minor differences to the scheduling problems considered so that the proof does not carry over formally. While the following complexity result is not unexpected, we formally include them for reasons of completeness. Moreover, it serves as a basis for the following design tasks in Sections III-B and III-C, which are tailored to ETCS L2 HTD.
@@
TR-VER is NP-complete even if all routes and exact stopping times in train stations are fixed and the railway network is planar.
@@
Given a certificate, we can check feasibility using the definitions in polynomial time; hence, TR-VER is in NP. To show NP-completeness, we reduce Monotone 3-SAT-(2,2$\leq$ 2$\leq$ 2≤ 2 , ≤ 2) to TR-VER. Monotone 3-SAT-(2,2$\leq$ 2$\leq$ 2≤ 2 , ≤ 2) are all SAT instances in which every clause contains three literals (all of which contain only unnegated or only negated variables), and every variable appears in clauses at most twice unnegated and twice negated. Monotone 3-SAT-(2,2$\leq$ 2$\leq$ 2≤ 2 , ≤ 2) is known to be NP-complete [45].
@@
Figure 14: Reduction from Monotone 3-SAT-(2,2$\leq$ 2$\leq$ 2≤ 2 , ≤ 2) to TR-VER